OpenStudy (anonymous):

Use the Midpoint Formula to find an endpoint C, such that B is the midpoint of AC. A(-2,1) B(1,-3). I really need help.

OpenStudy (anonymous):

C=-2,-0.5)

OpenStudy (anonymous):

How would I show my work I kept getting 3,-4 as my answer. Could you explain that? Please.

OpenStudy (anonymous):

you would do -3*1/2=-2 for one equation and 1*(-2)/2=-0.5 for the other equation

OpenStudy (anonymous):

That answer is not align with the slope.

OpenStudy (anonymous):

Not on the slope

OpenStudy (anonymous):



OpenStudy (anonymous):

i know i did it correct

OpenStudy (anonymous):

Where would it be plotted at because it mandatory that it is.